OpenStudy (anonymous):

Find all the zeros of the polynomial function f(x) = 2x3 - 8x2 - 6x + 24, where -3 is a root.

OpenStudy (jhannybean):

\[\large f(x)=2x^3-8x^2-6x+24\] factor by grouping.\[f(x)=2x^2(x-4)-6(x+4)\]regroup.\[f(x)=(2x^2-6)(x-4)\] equal equation to 0. and solve for x. \[(2x^2-6)=0\]\[(x-4)=0\]\[2x^2=6\]\[x^2=3\] take the square root of 3 and you will get two answers.\[x-4+4= 0+4\]\(x=4\) is another answer. solve the other one though.

OpenStudy (jhannybean):

you will have 3 answers. 2 positive, one negative.

OpenStudy (jhannybean):

thanks!